OpenStudy (anonymous):

What does \(\sum_{n\in\mathbb N}\frac{n^x}{x^n}\) converge to when \(x\in\mathbb R^+\) and \(x>1\)?

OpenStudy (anonymous):

\(\sum_{n\in\mathbb N}\frac{n^x}{x^n}\)

OpenStudy (anonymous):

Yes, I forget this isn't M.SE. :)

OpenStudy (anonymous):

just so i can read it, i don't know it

OpenStudy (anonymous):

i think if you use \[ instead of \( it will work

OpenStudy (anonymous):

I fixed it, refresh. The "\(" is for inline equations.

OpenStudy (anonymous):

aah now i can see it dollar signs don't work here for some reason

OpenStudy (anonymous):

I'm just trying to make sure this question is M.SE material before I post it, aha. Maybe it has a trivial solution I'm missing.

OpenStudy (anonymous):

is it supposed to be something nice?

OpenStudy (anonymous):

Not sure.

OpenStudy (jhannybean):

It's so tiny I can'r even read it.

OpenStudy (jhannybean):

\[\huge \sum_{n\in\mathbb N}\frac{n^x}{x^n}\]

OpenStudy (jhannybean):

@satellite73 I'm not sure what \(\large n \in \mathbb N\) means :|

OpenStudy (anonymous):

\(n\) is natural. \(x\) is positive reals.

OpenStudy (jhannybean):

So basically to infinity....yes ?

OpenStudy (anonymous):

Yes.

OpenStudy (jhannybean):

Ohh ok. Just to make sure... \[\large \sum_{n=1}^{\infty} \frac{n^x}{x^n}\]Yeah im not sure if n would start at 0 or1....

OpenStudy (jhannybean):

@mathstudent55 ? :\

OpenStudy (anonymous):

n starts at 1.

OpenStudy (dan815):

n E N just means n is all natural numbers

OpenStudy (dan815):